# Next Generation Equation-Free Multiscale Modelling of Crowd Dynamics via Machine Learning

### v2

@@ Appendix C Density fields extraction via Kernel Density Estimation
$\rho((x_{i},y_{j}),t)=\mathcal{M}(x_{i},y_{j})\,\rho((x_{i},y_{j}),t),\quad\forall(x_{i},y_{j})\in\Omega.$ (C.6)
@@ Appendix D Lag window size selection via the Bayesian and the Akaike Information Criteria
Appendix E Comparison of the “ground truth” density fields with the predicted ones, using MVAR(4), LSTM(4) and LSTM(9) models
@@
(b) Predicted density at $t_{250}=62.5\,\text{s}$.
(c) Ground truth density at $t_{500}=125\,\text{s}$.
(d) Predicted density at $t_{500}=125\,\text{s}$.
(e) Ground truth density at $t_{780}=195\,\text{s}$.
(f) Predicted density at $t_{780}=195\,\text{s}$.
(g) Ground truth density at time $t_{930}=232.5\,\text{s}$.
(h) Predicted density at $t_{930}=232.5\,\text{s}$.
@@
Figure E.1: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ normalized densities via the closed-loop time-stepper in Eq. (36) using the MVAR(4) model for an unseen case of the testing set; initialization at the 6th row of Table B.2. Panels (a), (c), (e), and (g) show the “ground truth” density distribution in $\Omega$, while panels (b), (d), (f), and (h) show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
(a) Ground truth density at $t_{250}=62.5\,\text{s}$.
(b) Predicted density at $t_{250}=62.5\,\text{s}$.
(c) Ground truth density at $t_{500}=125\,\text{s}$.
(d) Predicted density at $t_{500}=125\,\text{s}$.
(e) Ground truth density at $t_{780}=195\,\text{s}$.
(f) Predicted density at $t_{780}=195\,\text{s}$.
(g) Ground truth density at time $t_{930}=232.5\,\text{s}$.
(h) Predicted density at $t_{930}=232.5\,\text{s}$.
@@
Figure E.2: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ normalized densities via the closed-loop time-stepper in Eq. (36) using the LSTM(4) model for an unseen case of the testing set; initialization at the 6th row of Table B.2. Panels (a), (c), (e), and (g) show the “ground truth” density distribution in $\Omega$, while panels (b), (d), (f), and (h), show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
(a) Ground truth density at $t_{250}=62.5\,\text{s}$.
(b) Predicted density at $t_{250}=62.5\,\text{s}$.
(c) Ground truth density at $t_{500}=125\,\text{s}$.
(d) Predicted density at $t_{500}=125\,\text{s}$.
(e) Ground truth density at $t_{780}=195\,\text{s}$.
(f) Predicted density at $t_{780}=195\,\text{s}$.
(g) Ground truth density at time $t_{930}=232.5\,\text{s}$.
(h) Predicted density at $t_{930}=232.5\,\text{s}$.

### v3

@@ Appendix B Density fields extraction via Kernel Density Estimation
$\rho((x_{i},y_{j}),t)=\mathcal{M}(x_{i},y_{j})\,\rho((x_{i},y_{j}),t),\quad\forall(x_{i},y_{j})\in\Omega.$ (B.6)
Appendix C Initial Conditions of Microscopic Distributions
    $p(x_{i},y_{i})=\mathcal{U}\left([x_{\text{min}},x_{\text{max}}]\times[y_{\text{min}},y_{\text{max}}]\right).$ (C.1)
    $p(x_{i},y_{i})=\frac{1}{2\pi\sigma_{x}\sigma_{y}}\exp\left(-\frac{(x_{i}-\mu_{x})^{2}}{2\sigma_{x}^{2}}-\frac{(y_{i}-\mu_{y})^{2}}{2\sigma_{y}^{2}}\right),$ (C.2)
    $p(x_{i},y_{i})$ $=\frac{1}{\sqrt{2\pi}\,\sigma_{y}}\,\exp\!\left(-\frac{(y-\mu_{y})^{2}}{2\sigma_{y}^{2}}\right)$ (C.3)
    $p(x_{i},y_{i})=\frac{1}{\pi s_{x}s_{y}}\cos\left(\frac{x_{i}-c_{x}}{s_{x}}\right)\cos\left(\frac{y_{i}-c_{y}}{s_{y}}\right).$ (C.4)
The training and testing data sets were generated from $C=20$ initial conditions, with 5 samples drawn from each of the four distributions described above. To ensure plausible and varied crowd configurations, the parameters for these distributions were selected heuristically to provide a broad spatial coverage while mitigating boundary artifacts. Furthermore, during the sampling process, pedestrian positions were resampled if they were within a critical distance equal to $2r_{i}$ of an existing pedestrian or a wall/obstacle to prevent initial overlaps and collisions. The specific parameter configurations used for the training and testing data sets are detailed Table C.1 and Table C.2, respectively.
Table C.1: Microscopic initial distribution configurations and parameters for constructing the training dataset.
Case c Distribution Type Parameters
1 Uniform, Eq. (C.1) $x_{\text{min}}=2,x_{\text{max}}=15,y_{\text{min}}=3,y_{\text{max}}=9$
2 Uniform, Eq. (C.1) $x_{\text{min}}=33,x_{\text{max}}=46,y_{\text{min}}=3,y_{\text{max}}=9$
3 Uniform, Eq. (C.1) $x_{\text{min}}=2,x_{\text{max}}=20,y_{\text{min}}=4,y_{\text{max}}=8$
4 Uniform, Eq. (C.1) $x_{\text{min}}=28,x_{\text{max}}=46,y_{\text{max}}=4,y_{\text{max}}=8$
5 Uniform, Eq. (C.1) $x_{\text{min}}=3,x_{\text{max}}=17,y_{\text{max}}=4,y_{\text{max}}=8$
6 Gaussian, Eq. (C.2) $\mu_{x}=10,\;\mu_{y}=6,\sigma_{x}=1.5,\;\sigma_{y}=1.5$
7 Gaussian, Eq. (C.2) $\mu_{x}=10,\;\mu_{y}=6,\sigma_{x}=2,\;\sigma_{y}=1.5$
8 Gaussian, Eq. (C.2) $\mu_{x}=38,\;\mu_{y}=6,\sigma_{x}=2,\;\sigma_{y}=1.5$
9 Gaussian, Eq. (C.2) $\mu_{x}=38,\;\mu_{y}=6,\sigma_{x}=1.5,\;\sigma_{y}=1.5$
10 Gaussian, Eq. (C.2) $\mu_{x}=37,\;\mu_{y}=5,\sigma_{x}=1.5,\;\sigma_{y}=1.5$
11 Double Gaussian, Eq. (C.3) $\mu_{x_{1}}=8,\;\mu_{y}=7,\sigma_{x_{1}}=1.2,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=12,\;\sigma_{x_{2}}=0.8$
12 Double Gaussian, Eq. (C.3) $\mu_{x_{1}}=10,\;\mu_{y}=6,\sigma_{x_{1}}=1.2,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=14,\;\sigma_{x_{2}}=1$
13 Double Gaussian, Eq. (C.3) $\mu_{x_{1}}=28,\;\mu_{y}=6,\sigma_{x_{1}}=1.3,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=32,\;\sigma_{x_{2}}=1.1$
14 Double Gaussian, Eq. (C.3) $\mu_{x_{1}}=36,\;\mu_{y}=6,\sigma_{x_{1}}=1.4,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=40,\;\sigma_{x_{2}}=1.2$
15 Double Gaussian, Eq. (C.3) $\mu_{x_{1}}=9,\;\mu_{y}=6,\sigma_{x_{1}}=0.8,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=11,\;\sigma_{x_{2}}=1$
16 Cosine, Eq. (C.4) $c_{x}=10,\;c_{y}=6,s_{x}=10,\;s_{y}=3$
17 Cosine, Eq. (C.4) $c_{x}=20,\;c_{y}=8,s_{x}=10,\;s_{y}=3$
18 Cosine, Eq. (C.4) $c_{x}=30,\;c_{y}=7,s_{x}=10,\;s_{y}=3$
19 Cosine, Eq. (C.4) $c_{x}=40,\;c_{y}=6,s_{x}=10,\;s_{y}=3$
20 Cosine, Eq. (C.4) $c_{x}=10,\;c_{y}=6,s_{x}=9,\;s_{y}=3$
Table C.2: Microscopic initial distribution configurations and parameters for constructing the testing dataset.
Case c Distribution Type Parameters
1 Uniform, Eq. (C.1) $x_{\text{min}}=2,\;x_{\text{max}}=16,\;y_{\text{min}}=3,\;y_{\text{max}}=10$
2 Uniform, Eq. (C.1) $x_{\text{min}}=32,\;x_{\text{max}}=46,\;y_{\text{min}}=3,\;y_{\text{max}}=10$
3 Uniform, Eq. (C.1) $x_{\text{min}}=2,\;x_{\text{max}}=19,\;y_{\text{min}}=4,\;y_{\text{max}}=8$
4 Uniform, Eq. (C.1) $x_{\text{min}}=29,\;x_{\text{max}}=46,\;y_{\text{min}}=4,\;y_{\text{max}}=8$
5 Uniform, Eq. (C.1) $x_{\text{min}}=4,\;x_{\text{max}}=17,\;y_{\text{min}}=4,\;y_{\text{max}}=8$
6 Gaussian, Eq. (C.2) $\mu_{x}=12,\;\mu_{y}=7,\;\sigma_{x}=2,\;\sigma_{y}=1.5$
7 Gaussian, Eq. (C.2) $\mu_{x}=11,\;\mu_{y}=5,\;\sigma_{x}=1.5,\;\sigma_{y}=2$
8 Gaussian, Eq. (C.2) $\mu_{x}=38,\;\mu_{y}=6,\;\sigma_{x}=2,\;\sigma_{y}=1.5$
9 Gaussian, Eq. (C.2) $\mu_{x}=39,\;\mu_{y}=5,\;\sigma_{x}=1.5,\;\sigma_{y}=1.5$
10 Gaussian, Eq. (C.2) $\mu_{x}=37,\;\mu_{y}=7,\;\sigma_{x}=1.5,\;\sigma_{y}=1.5$
11 Double Gaussian, Eq. (C.3) $\mu_{x_{1}}=7,\;\mu_{y}=7,\;\sigma_{x_{1}}=1.1,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=11,\;\sigma_{x_{2}}=0.9$
12 Double Gaussian, Eq. (C.3) $\mu_{x_{1}}=10,\;\mu_{y}=7,\;\sigma_{x_{1}}=1.2,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=14,\;\sigma_{x_{2}}=1$
13 Double Gaussian, Eq. (C.3) $\mu_{x_{1}}=29,\;\mu_{y}=6,\;\sigma_{x_{1}}=1.3,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=33,\;\sigma_{x_{2}}=1.1$
14 Double Gaussian, Eq. (C.3) $\mu_{x_{1}}=37,\;\mu_{y}=6,\;\sigma_{x_{1}}=1.4,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=41,\;\sigma_{x_{2}}=1.2$
15 Double Gaussian, Eq. (C.3) $\mu_{x_{1}}=10,\;\mu_{y}=6,\;\sigma_{x_{1}}=0.8,\;\sigma_{y}=1.5,\;\mu_{x_{2}}=12,\;\sigma_{x_{2}}=1$
16 Cosine, Eq. (C.4) $c_{x}=15,\;c_{y}=4,\;s_{x}=8,\;s_{y}=3$
17 Cosine, Eq. (C.4) $c_{x}=25,\;c_{y}=6,\;s_{x}=10,\;s_{y}=3$
18 Cosine, Eq. (C.4) $c_{x}=35,\;c_{y}=5,\;s_{x}=10,\;s_{y}=3$
19 Cosine, Eq. (C.4) $c_{x}=45,\;c_{y}=7,\;s_{x}=5,\;s_{y}=3$
20 Cosine, Eq. (C.4) $c_{x}=15,\;c_{y}=6,\;s_{x}=7,\;s_{y}=3$
@@ Appendix D Lag window size selection via the Bayesian and the Akaike Information Criteria
Appendix E Open-loop prediction errors and closed-loop reconstructions for the unidirectional flow case
Here, we provide the detailed open-loop prediction errors and the closed-loop reconstructions of the proposed framework for the unidirectional flow case discussed in Section 4.1.
The open-loop setting prediction of the proposed restrict-evolve with ROM-lift framework in Eq. (5) reads:
$\hat{x}^{(c)}(t_{k})=\mathcal{L}\left(\hat{\boldsymbol{y}}^{(c)}(t_{k})\right),\qquad\hat{\boldsymbol{y}}^{(c)}(t_{k})=\Phi_{\text{ROM}}\left(\boldsymbol{y}^{(c)}(t_{k-1}),\ldots,\boldsymbol{y}^{(c)}(t_{k-w});\mathbf{p}\right),$ (E.1)
where $\boldsymbol{y}^{(c)}(t_{k-j})=\mathcal{R}(x^{(c)}(t_{k-j}))$ are restricted embeddings of observed densities for $j=1,\ldots,w$. Similarly to the close-loop predictions in Eq. (51), the open-loop prediction expression is used in Eq. (52) to quantify the one-step prediction accuracy of the proposed framework.
Table E.1 provides a summary of the relative reconstructed errors in the ambient space, for one-step predictions, for all four ROMs, including the mean error and the 10–90% error percentiles, aggregated over all cases and time steps in the testing set.
Table E.1: Open-loop (one-step) prediction errors in the ambient–density profile–space, for the testing set using the trained MVAR and LSTM (best out of 50 training runs) models for the unidirectional flow case. The relative reconstructed $L_{1}$, $L_{2}$ and $L_{\infty}$ errors $e^{1,(c)}_{k},\ e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ in Eq. (52) are reported for the MVAR(4), MVAR(9), LSTM(4), and LSTM(9) latent dynamics models. Mean and 10–90% error percentiles are shown over all the $C=20$ cases and $K=\{991,996\}$ time steps (for width $w=\{9,4\}$, respectively).
MVAR(4) 0.048(0.033, 0.065)0.048\ (0.033,\,0.065) 0.038(0.027, 0.051)0.038\ (0.027,\,0.051) 0.042(0.030, 0.057)0.042\ (0.030,\,0.057)
MVAR(9) 0.047(0.032, 0.063)0.047\ (0.032,\,0.063) 0.037(0.026, 0.049)0.037\ (0.026,\,0.049) 0.041(0.030, 0.055)0.041\ (0.030,\,0.055)
LSTM(4) 0.048(0.034, 0.065)0.048\ (0.034,\,0.065) 0.038(0.027, 0.051)0.038\ (0.027,\,0.051) 0.043(0.032, 0.058)0.043\ (0.032,\,0.058)
LSTM(9) 0.047(0.033, 0.063)0.047\ (0.033,\,0.063) 0.037(0.026, 0.049)0.037\ (0.026,\,0.049) 0.042(0.031, 0.056)0.042\ (0.031,\,0.056)
Figure E.1 displays the relative reconstruction errors $e^{2,(c)}_{k},\ e^{\infty,(c)}_{k}$ and their 10–90% percentile ranges for one-step predictions via the open-loop time-stepper (Eq. (E.1)). Errors are averaged over the $C=20$ cases of different, unseen initial conditions considered, across the time steps $k=1,\ldots,1000$.
(a) Relative Errors with MVAR(4)
(b) Relative Errors with MVAR(9)
(c) Relative Errors with LSTM(4)
(d) Relative Errors with LSTM(9)
Figure E.1: One-step (open-loop) prediction errors, in the ambient–density profile–space, for the testing set over time, using trained MVAR and LSTM models for the unidirectional flow case. The relative reconstructed $L_{2}$ (cyan) and $L_{\infty}$ (green) errors $e^{2,(c)}_{k}$ and $e^{\infty,(c)}_{k}$ Eq. (52) (using the one-step prediction in Eq. (E.1)) are shown for the MVAR(4) panel (a), MVAR(9) panel (b), LSTM(4) panel (c), and LSTM(9) panel (d) latent dynamics models. Mean relative errors (solid) and 10–90% error percentiles (dashed) are shown over $C=20$ cases per time step. For the LSTM models, results correspond to the model achieving the best training performance out of the 50 independent runs.
Figures E.2, E.3 E.4 and E.5 compare the ground truth normalized density fields $x^{(c)}(t_{k})$ and the predicted ones $\hat{x}^{(c)}(t_{k})$ for closed-loop predictions (Eq. (51)) using MVAR(4), MVAR(9), LSTM(4), and LSTM(9) models, respectively. The test case uses a Gaussian initialization (6th row of Table C.2) at four time steps ($k=250,500,780,930$), during which the pedestrian crowd does not pass through the corridor boundary. Panels (a,c,e,g) show the ground truth, while panels (b,d,f,h) present the corresponding closed-loop predictions. The spatial distribution of the resulting absolute errors is shown in Figure E.6 at time steps $k=250,930$.
(a) Ground truth density at $t_{250}=62.5\,\text{s}$
(b) Predicted density at $t_{250}=62.5\,\text{s}$
(c) Ground truth density at $t_{500}=125\,\text{s}$
(d) Predicted density at $t_{500}=125\,\text{s}$
(e) Ground truth density at $t_{780}=195\,\text{s}$
(f) Predicted density at $t_{780}=195\,\text{s}$
(g) Ground truth density at $t_{930}=232.5\,\text{s}$
(h) Predicted density at $t_{930}=232.5\,\text{s}$
Figure E.2: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ normalized densities via the closed-loop time-stepper in Eq. (51) for the unidirectional flow case, using the MVAR(4) model for an unseen case of the testing set; initialization at the 6th row of Table C.2. Panels (a), (c), (e), and (g) show the “ground truth” density distribution in $\Omega$, while panels (b), (d), (f), and (h) show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
@@ Appendix E Open-loop prediction errors and closed-loop reconstructions for the unidirectional flow case
(b) Predicted density at $t_{250}=62.5\,\text{s}$
(c) Ground truth density at $t_{500}=125\,\text{s}$
(d) Predicted density at $t_{500}=125\,\text{s}$
(e) Ground truth density at $t_{780}=195\,\text{s}$
(f) Predicted density at $t_{780}=195\,\text{s}$
(g) Ground truth density at $t_{930}=232.5\,\text{s}$
(h) Predicted density at $t_{930}=232.5\,\text{s}$
@@
Figure E.3: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ normalized densities via the closed-loop time-stepper in Eq. (51) for the unidirectional flow case, using the MVAR(9) model for an unseen case of the testing set; initialization at the 6th row of Table C.2. Panels (a), (c), (e), and (g) show the “ground truth” density distribution in $\Omega$, while panels (b), (d), (f), and (h), show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
(a) Ground truth density at $t_{250}=62.5\,\text{s}$
(b) Predicted density at $t_{250}=62.5\,\text{s}$
(c) Ground truth density at $t_{500}=125\,\text{s}$
(d) Predicted density at $t_{500}=125\,\text{s}$
(e) Ground truth density at $t_{780}=195\,\text{s}$
(f) Predicted density at $t_{780}=195\,\text{s}$
(g) Ground truth density at $t_{930}=232.5\,\text{s}$
(h) Predicted density at $t_{930}=232.5\,\text{s}$
@@
Figure E.4: Comparison of the “ground truth” $x^{(c)}(t_{k})$ and the predicted $\hat{x}^{(c)}(t_{k})$ normalized densities via the closed-loop time-stepper in Eq. (51) for the unidirectional flow case, using the LSTM(4) model for an unseen case of the testing set; initialization at the 6th row of Table C.2. Panels (a), (c), (e), and (g) show the “ground truth” density distribution in $\Omega$, while panels (b), (d), (f), and (h), show the predicted one, for the time steps $t_{250}=62.5\,\text{s}$, $t_{500}=125\,\text{s}$, $t_{780}=195\,\text{s}$, and $t_{930}=232.5\,\text{s}$, respectively.
(a) Ground truth density at $t_{250}=62.5\,\text{s}$
(b) Predicted density at $t_{250}=62.5\,\text{s}$
(c) Ground truth density at $t_{500}=125\,\text{s}$
(d) Predicted density at $t_{500}=125\,\text{s}$
(e) Ground truth density at $t_{780}=195\,\text{s}$
(f) Predicted density at $t_{780}=195\,\text{s}$
(g) Ground truth density at $t_{930}=232.5\,\text{s}$
(h) Predicted density at $t_{930}=232.5\,\text{s}$